find the solution to the system of equations.
y = 3x -3
y = 5x - 4

Answer by checkley77(12844)

y = 3x -3
y = 5x - 4
3X-3=5X-4
3X-5X=-4+3
-2X=-1
X=1/2
X=1/2 OR .5
Y=3*.5-3
Y=1.5-3
Y=-1.5 ANS
PROOF:
-1.5=5*1/2-4
-1.5=2.5-4
-1.5=-1.5